OpenStudy (anonymous):

A body is travelling with a constant acceleration of 10m/s2.If S1 is the dsitance travelled in 1st sec and S2 is the ditance travelled in 2nd sec which of the follwing show correct relation between between S1 and S2? A-S1=S2 B-S1=3S2 C-S2=3S1 D-none

OpenStudy (anonymous):

@Imtiaz7 @iambatman @myininaya @whpalmer4 @robtobey

OpenStudy (anonymous):

@random231 @ShadowLegendX

OpenStudy (anonymous):

I think it same because Acceleration is same

random231 (random231):

no

OpenStudy (anonymous):

:/

random231 (random231):

find the distance traveled in first second using s = ut +1/2 * at^2

OpenStudy (anonymous):

value of u ??

OpenStudy (anonymous):

what is value of a then ?

random231 (random231):

oh shoot my bad

random231 (random231):

take initial vel as u and a= 10 fid the distance in terms of u

OpenStudy (kropot72):

The average speed in the first second is: \[\frac{0+10}{2}=5\ m/s\] The average speed in the second second is: \[\frac{10+20}{2}=15\ m/s\] So can you now see the ratio of the distance traveled in the first second to the distance traveled in the 2nd second.

OpenStudy (anonymous):

Got it . Its C Thank you @kropot72 ^_^

OpenStudy (kropot72):

You're welcome :)